# TDFlow: Agentic Workflows for Test Driven Software Engineering

### v1

@@ Abstract
TDFlow: Agentic Workflows for Test Driven Software Engineering
@@ 2.2 Software Engineering Agents and Benchmarks
Most repository scale software engineering agent systems rely on single ReACT agents, or multiple agents, equipped with a set of tools and tasked with solving the issue (Wang et al., 2024d; Yang et al., 2024a; Tao et al., 2024; Wang et al., 2024c, b; Yao et al., 2023). Some examples of this include OpenDevin/OpenHands, the first of such work, and SWE-agent (Wang et al., 2024c, d; Yang et al., 2024a). Multi-agent systems involve multiple autonomous agents each given personas such as "Manager", "Fault Localizer", and "Verifier" and tasked with different objectives. Such systems include MAGIS, CODER, and AgentCoder (Tao et al., 2024; Chen et al., 2024; Huang et al., 2024).
@@
Automated test generation systems such as CodeT, Libro, Otter, UTBoost and AEGIS also utilize LLMs in both workflow and agentic fashions for generating reproduction tests of repository scale issues (Chen et al., 2023; Kang et al., 2023; Wang et al., 2024a; Ahmed et al., 2025). Benchmarks such as TDD-Bench and SWT-Bench also provide a repository scale benchmark on the quality of LLM-generated reproduction tests (Mündler et al., 2024; Ahmed et al., 2024). As the focus of our work lies in resolving human-written tests, we consider other automated test generation systems and the evaluation thereof as complementary work.
@@ 4.2 The final hurdle to human-level accuracy lies in test generation
The results of the experiment are found in Table 2. When provided with human-written tests, TDFlow achieves a 94.3% success rate on SWE-Bench Verified, approaching human-level performance on test resolution. Combined with the 68.0% performance when generating its own tests, this suggests the test resolution capabilities of TDFlow are already human-level, while the final hurdle in LLM respository-scale issue resolution lies within test generation. Furthermore, the test resolution cost in the LLM-generated mode is significantly greater, most likely due to generated tests which don’t truly test reproduction behavior.
@@ 7 Conclusion
In this work, we present TDFlow, a test-driven agentic workflow specifically designed to solve human-written reproduction tests in a test-driven setting. TDFlow is a simple, precisely engineered, modular workflow consisting of several sub-agents each tasked with a well-defined subtask and provided the minimum set of tools and limited agentic freedom. The modularity of TDFlow allows for the decoupling and individual, targeted training of each sub-agent on specifically-designed tasks. We benchmark TDFlow on SWE-Bench Lite with GPT-4.1 and find it outperforms the next best system by 27.8% when provided human-written tests. On SWE-Bench Verified, when solving successful reproduction tests, TDFlow already shows human-level performance of 94.3% and 93.3% pass rate regardless if the test were human-written or LLM-generated. Finally, we show that the final obstacle to autonomous software engineering, in the context of test-driven development and the SWE-Bench benchmark, lies in generating better tests.
@@ 8 Limitations
By nature of TDFlow being a rigid workflow consisting of multiple sub-agents instead of a single adaptable agent, the environment, infrastructure, and test information setup is more complex and less flexible than purely agentic systems. Furthermore, if the provided tests are truly unsolvable, TDFlow will keep performing main algorithm iterations until all of the iterations have been exhausted. There is no early-stopping mechanism or critic that can be used to save resources.

### v2

# TDFlow: Agentic Workflows for Test Driven Development

@@ Abstract
TDFlow: Agentic Workflows for Test Driven Development
@@ 2.2 Software Engineering Agents and Benchmarks
Most repository scale software engineering agent systems rely on single ReACT agents, or multiple agents, equipped with a set of tools and tasked with solving the issue (Wang et al., 2024c; Yang et al., 2024a; Tao et al., 2024; Wang et al., 2024b, a; Yao et al., 2023). Some examples of this include OpenDevin/OpenHands, the first of such work, and SWE-agent (Wang et al., 2024b, c; Yang et al., 2024a). Multi-agent systems involve multiple autonomous agents each given personas such as "Manager", "Fault Localizer", and "Verifier" and tasked with different objectives. Such systems include MAGIS, CODER, and AgentCoder (Tao et al., 2024; Chen et al., 2024; Huang et al., 2024).
@@
Automated test generation systems such as CodeT, Libro, Otter, UTBoost and AEGIS also utilize LLMs in both workflow and agentic fashions for generating reproduction tests of repository scale issues (Chen et al., 2023; Kang et al., 2023; Wang and others, 2024; Ahmed et al., 2025). Benchmarks such as TDD-Bench and SWT-Bench also provide a repository scale benchmark on the quality of LLM-generated reproduction tests (Mündler et al., 2024; Ahmed et al., 2024). As the focus of our work lies in resolving human-written tests, we consider other automated test generation systems and the evaluation thereof as complementary work.
@@ 4.2 The final hurdle to human-level accuracy lies in test generation
The results of the experiment are found in Table 2. When provided with human-written tests, TDFlow achieves a 94.3% success rate on SWE-Bench Verified, approaching human-level performance on test resolution. Combined with the 68.0% performance when generating its own tests, this suggests the test resolution capabilities of TDFlow are already human-level, while the final hurdle in LLM repository-scale issue resolution lies within test generation. Furthermore, the test resolution cost in the LLM-generated mode is significantly greater, most likely due to generated tests which don’t truly test reproduction behavior.
@@ 7 Conclusion
In this work, we present TDFlow, a test-driven agentic workflow specifically designed to solve human-written reproduction tests in a test-driven setting. TDFlow is a simple, precisely engineered, modular workflow consisting of several sub-agents each tasked with a well-defined subtask and provided the minimum set of tools and limited agentic freedom. The modularity of TDFlow allows for the decoupling and individual, targeted training of each sub-agent on specifically-designed tasks. We benchmark TDFlow on SWE-Bench Lite with GPT-4.1 and find it outperforms the next best system by 27.8% when provided human-written tests. On SWE-Bench Verified, when solving successful reproduction tests, TDFlow already shows human-level performance of 94.3%. Finally, through an analysis of the TDFlow system, we show that the final obstacle to autonomous software engineering, in the context of test-driven development and the SWE-Bench benchmark, lies in generating better reproduction tests.
@@ 8 Limitations
By nature of TDFlow being a rigid workflow consisting of multiple sub-agents instead of a single adaptable agent, the environment, infrastructure, and test information setup is more complex and less flexible than purely agentic systems. Furthermore, if the provided tests are truly unsolvable, TDFlow will keep performing main algorithm iterations until the iteration limit has been reached. There is no early-stopping mechanism or critic that can be used to save resources.